Write the next four terms of the sequence defined recursively.
a1=14
an+1=(-2)an

Answer by greenestamps(13216)



The recursive definition

a(n+1) = (-2)*a(n)

means that each term is found by multiplying the previous term by -2.

So the definition of this sequence says that the first term is 14, and each of the following terms is the preceding term, multiplied by -2:
a(1) = 14
a(2) = -2(14) = -28
a(3) = -2(-28) = 56
a(4) = -2(56) =-112